OpenStudy (anonymous):

Find the two digit number whose tens digit is 4 less than its units digit if the original number is 2 more than 3 times the sum of the digits.

OpenStudy (anonymous):

Let u be the magnitude of the units digit. The following equation can be derived from the problem statement: 10 (u - 4) + u = 2 + 3 (u - 4 + u) u = 6 The number is 26. 26 = 2 + 3(6 + 2)